OpenStudy (anonymous):

12/13x + 1/2 = 1/3 - 1/13x + 1/3 I am so confused on how to solve this.. help anyone? by the way the / is for a fraction not division..

OpenStudy (anonymous):

\[\frac{ 12 }{ 13x } + \frac{ 1 }{ 2 }= \frac{ 1 }{ 3 }- \frac{ 1 }{ 13x } +\frac{ 1 }{3}\]

OpenStudy (anonymous):

find the lcm and then simplify

OpenStudy (anonymous):

Okay for the left side, would the lcm be 26? Then for the right..39? I'm so confused! haha. Or would it be.. 2 and 3?

OpenStudy (anonymous):

the lcm is (2)(3)(13x)

OpenStudy (anonymous):

ahh, okay. So do we combine the like terms now? l:

OpenStudy (ageta):

Answer: x=1/6

OpenStudy (anonymous):

yes... combine it and tell me the answer u got

OpenStudy (anonymous):

\[\frac{ 6(12)+3(13x)=2(13x)-6+2(13x) }{ (2)(3)(13x) }\]

OpenStudy (anonymous):

now you can consider the numerator...do the multiplication and then collect like terms and find x

Directrix (directrix):

Do we know for sure on this problem that the x variables are in the denominator? I'm wondering if the equation should be: (12/13) x + 1/2 = 1/3 - (1/13) x + 1/3 @kaela_stanley

OpenStudy (anonymous):

You're right... @Directrix I didn't even notice that... -_-